# Machine Learning Framework for Characterizing Processing–Structure Relationship in Block Copolymer Thin Films

### v1

@@ Abstract
The morphology of block copolymers (BCPs) critically influences material properties and applications. This work introduces a machine learning (ML)-enabled, high-throughput framework for analyzing grazing incidence small-angle X-ray scattering (GISAXS) data and atomic force microscopy (AFM) images to characterize BCP thin film morphology. A convolutional neural network was trained to classify AFM images by morphology type, achieving 97% testing accuracy. Classified images were then analyzed to extract 2D grain size measurements from the samples in a high-throughput manner. ML models were developed to predict morphological features based on processing parameters such as solvent ratio, additive type, and additive ratio. GISAXS-based properties were predicted with strong performances ($R^{2}$ > 0.75), while AFM-based property predictions were less accurate ($R^{2}$ < 0.60), likely due to the localized nature of AFM measurements compared to the bulk information captured by GISAXS. Beyond model performance, interpretability was addressed using Shapley Additive exPlanations (SHAP). SHAP analysis revealed that the additive ratio had the largest impact on morphological predictions, where additive provides the BCP chains with increased volume to rearrange into thermodynamically favorable morphologies. This interpretability helps validate model predictions and offers insight into parameter importance. Altogether, the presented framework combining high-throughput characterization and interpretable ML offers an approach to exploring and optimizing BCP thin film morphology across a broad processing landscape.
@@
[…]
@@ 1 Introduction
Copolymer thin films consist of a wide range of materials with a vast array of applications spanning from membrane technologies for gas separation to energy applications including solid electrolytes.1, 2, 3 They exhibit a diverse range of tunable properties dependent on both chemical structure and morphology. Copolymers, specifically block copolymers (BCPs), are a class of polymers in which multiple chemically distinct polymer chains, as building blocks, are bonded together forming one macromolecule. When the blocks of a copolymer are immiscible, unfavorable interactions between the blocks lead to phase separation resulting in unique morphologies.4 These unique nanoscale phase-separated microdomains can be leveraged to facilitate the design of advanced materials.5, 6 For example, BCP membranes find higher applicability compared to traditional inorganic membrane materials in water purification, battery electrolytes, and gas separation fields.7
@@
BCP self-assembly behavior at thermodynamic equilibrium can be accurately predicted using self-consistent field theory.8, 9, 10 The theory-predicted morphology of BCPs is primarily dictated by interactions between the blocks, known as the Flory-Huggins interaction parameter ${\chi}$, the degree of polymerization $N$, and the volume fraction of each block ${\phi}$. 11 In solid state, the phase diagram can be more complex due to polymer chains becoming kinetically trapped during sample preparation, resulting in a non-equilibrium state. Additionally, interfacial energies and confinement effects further increase the complexity of BCP self-assembly behavior.12, 13, 14
@@
To achieve precise control over BCP solid state morphology, processing conditions such as solvent type, sample casting method, and post-deposition annealing technique require optimization.15 The importance of solvent type is especially pronounced when the solvent is selective toward a specific block. In this case, changing the selectivity of the solvent may result in changes in the BCP morphology.16 Additionally, annealing techniques such as thermal and solvent-vapor annealing of BCP thin films play a large role in the final morphology. These techniques allow for increased mobility of the polymer chains, leading to a more thermodynamically favorable morphology.7 Moreover, casting methods such as drop casting and spin coating of BCP solution also affect the morphology mainly due to variation in solvent evaporation dynamics. Spin coating promotes rapid solvent removal and shear induced alignment resulting in uniform films with well-ordered microdomains and reduced defect densities.17, 18 On the other hand, slower solvent evaporation during drop casting leads to thicker films with varied domain orientation and rougher surface features.17, 18, 19 Therefore, these differences affect overall phase behavior and domain spacing, as the interplay between solvent evaporation, interfacial energetics, and film thickness determine the degree of phase separation and ordering.19, 20, 21
@@
Scattering and microscopy are common techniques for morphological characterization of BCP films. Grazing incidence small-angle X-ray scattering (GISAXS) is an inverse-space characterization technique measuring the phase behavior of BCP films to extract the average domain spacing and degree of ordering, where the domain of a BCP refers to the aggregation of a specific block.22 Atomic force microscopy (AFM) is a real-space characterization technique that can provide insights in BCP surface morphology and serve as a complementary characterization method to GISAXS.23 Moreover, periodically ordered arrangements of BCP domains, i.e., grains, can be identified and measured in AFM as well. However, analysis of characterization data of BCPs remains time-consuming, as manual analysis is often required, especially for features such as grain size from AFM images.
@@
Despite yielding great achievements, the traditional process of BCP solid state morphology control is constrained by laborious experimentation and data analysis required from researchers for the optimization of processing conditions. High-throughput data analysis and machine learning (ML) algorithms have demonstrated promising results in the field of materials science, offering effective approaches for solving optimization problems.24, 25, 26 However, these algorithms typically require large datasets for the training process. Due to the complexity of polymer systems, such as their intricate chemical structures, the availability of specific polymer datasets is limited, necessitating high-throughput experiments to generate the required data.27, 28, 29, 30 Therefore, high-throughput data acquisition, combined with ML-aided data analysis and interpretation, can be leveraged to explore the processing parameter space for controlling the solid-state morphology of BCPs.
@@
ML algorithms are extremely diverse, tailored for specific tasks and datasets. These algorithms seek to identify relationships within a labeled training dataset, in the case of supervised learning, optimizing the performance metric.31 Within supervised learning, there are two primary tasks: regression and classification. The key difference between these two tasks is the output: a numerical value for regression and a categorical output for classification.32 In the field of materials science, these models have provided researchers with an approach to reducing experimentation by predicting the properties of untested systems, as well as a tool to understand the relationships within complex material systems.33, 34, 35, 36 Numerous models have been applied in the materials science field to great success, ranging from random forest (RF), support vector machine (SVM), multilayer perceptron (MLP), and extreme gradient boosting (XGB).37, 38, 39, 40, 41, 42, 43, 44 Within the polymer field, ML has been leveraged to decipher scattering profiles and AFM images of polymers, elucidating polymer conformational and morphological features.45, 46, 47, 48 Specifically, the Jayaraman group has pioneered the integration of ML and molecular simulations to advance the interpretation of scattering data from complex polymer solutions. Their approach enables the analysis of scattering profiles that lack conventional fitting equations, overcoming a major limitation in characterizing novel polymer architectures. These works significantly broaden the applicability of scattering experiments, offering new insights into the structure and behavior of emerging polymer systems. 49, 50, 51
@@
Despite the widespread use of ML for polymer property predictions in the solution and solid state, comparatively little work has focused on the utilization of ML to understand how processing conditions influence the solid-state morphology of BCPs – an area critical for the design of functional materials. Furthermore, generating large datasets in the field of polymer science remains a significant challenge due to long experimentation and analysis times. Many studies have circumvented this issue by compiling pre-existing data from online databases and published papers into a single dataset.52, 53, 54, 24 However, this compiled data can be highly variable, as it comes from different laboratories, leading to inconsistencies in both the experimental setups and data analysis, as well as missed or mislabeled descriptors.
@@
Therefore, the synergistic use of high-throughput data analysis and ML methods can address the previously mentioned challenges and accelerate the rational design of polymeric materials.55 In this paper, we present an approach for the high-throughput data analysis of BCP thin film GISAXS profiles and two-dimensional grain sizes measured by AFM, creating a robust dataset of morphological properties from over 200 BCP thin films. In addition, interpretable ML is leveraged to evaluate the relationship between processing parameters and BCP morphology, and offers direct morphological property predictions from processing conditions. This work serves as a proof of concept for an integrated high-throughput data analysis-ML framework for the optimization of processing conditions for polymeric materials (Fig. 1).
@@ 2.1 Experimentation and characterization
Polystyrene-block-poly (ethylene oxide) (PS-b-PEO) (Mn = 23-b-7 kg/mol) was used as a model system to generate thin films using the “nonvolatile additive solvent annealing” (NVASA) approach. The details of the NVASA process can be found in our previous publication.56 PS-b-PEO was dissolved overnight in host solvents composed of mixtures of toluene and tetrahydrofuran (THF) at the following volume ratios, 100:0, 90:10, 80:20, 70:30, 60:40 and 50:50, to form 2 to 3 wtstock solutions. With NVASA, it has been previously reported that a controlled amount of high boiling point additive can briefly remain in the polymer films after the primary low boiling point solvent evaporates after spin coating, thereby swelling the BCP thin film, providing segmental mobility and improving domain order. 57 Therefore, for each stock solution, varied masses of a high boiling point additive: chloronaphthalene (CN), or methylnaphthalene (MN), were added to achieve additive ratios from 1.00 to 7.00. The control of additive ratio is dictated by the mass of additive relative to the mass of BCP (Equation 1).
Additive $\text{Additive Ratio}=\frac{mass_{additive}}{mass_{BCP}}+1$ (1)
@@
Oxford Cypher Asylum AFM was used to image the surface morphology of BCP thin films. Phase and height images (1 $\mu$m $\times$ 1 $\mu$m) with resolution of 256 $\times$ 256 pixels were taken in tapping mode. The tapping mode cantilevers (Tap300AI-G) with force constant of 40 N/m and resonant frequency of 300 kHz were used to scan the surface of films at a rate of 1.95 Hz. AFM raw data were collected using Igor Pro–based software, then flattened, and exported as images using Gwyddion software. These images were later processed in high-throughput fashion to extract properties of interest.
@@ 2.2.1 GISAXS
In this equation, $I(q)$ represents the scattering intensity as a function of the scattering vector $q$. The constant $A$ accounts for a baseline background offset, while the term $Cq^{-d}$ models a power law decay commonly associated with background or form factor scattering, with $C$ as the amplitude and $d$ as the decay exponent. The Gaussian peak is characterized by an amplitude $D$, a center position $b$, and a full width at half maximum (FWHM) $w$. The peak position $b$ is used to calculate the domain spacing $ds$, using the relation $ds=2\pi/b$. The FWHM, $w$, reflects the distribution of domain spacings and serves as an indicator of the degree of structural order in the sample. This fitting function enables simultaneous characterization of both periodic structural features and background scattering behavior.58 An in-house Python script was used to fit the scattering data, minimizing the sum of squared residuals within 50,000 maximum function evaluations. The script outputs the final fitted values for peak position and FWHM, along with the coefficient of determination ($R^{2}$) of the fitted function. This process for GISAXS analysis is both highly accurate and computationally efficient, with an average fitting $R^{2}$ of 0.99, taking less than 1 second per sample (Fig. 2).

### v2

@@ Abstract
The morphology of block copolymers (BCPs) critically influences material properties and applications. This work introduces a machine learning (ML)-enabled, high-throughput framework for analyzing grazing incidence small-angle X-ray scattering (GISAXS) data and atomic force microscopy (AFM) images to characterize BCP thin film morphology. A convolutional neural network was trained to classify AFM images by surface features, achieving 97% testing accuracy. Classified images were then analyzed to extract 2D grain size measurements from the samples in a high-throughput manner. ML models were trained to predict domain orientation based on processing parameters such as solvent ratio, additive type, and additive ratio. GISAXS-based properties were predicted with strong performances ($R^{2}$ > 0.75), while AFM-based property predictions were less accurate ($R^{2}$ < 0.60), likely due to the localized nature of AFM measurements compared to the bulk information captured by GISAXS. Beyond model performance, interpretability was addressed using Shapley Additive exPlanations (SHAP). SHAP analysis revealed that the additive ratio had the largest impact on morphological predictions, where additive provides the BCP chains with increased volume to rearrange into thermodynamically favorable morphologies. This interpretability helps validate model predictions and offers insight into parameter importance. Altogether, the presented framework combining high-throughput characterization and interpretable ML offers an approach to exploring and optimizing BCP thin film morphology across a broad processing landscape.
@@
These $\altaffiliationThese authors contributed equally to this work.\altaffiliationUniversity$ of Grenoble Alpes, CEA, Leti, Grenoble, F-38000 France
@@ 1 Introduction
Copolymer thin films consist of a wide range of materials with a vast array of applications spanning from membrane technologies for gas separation to energy applications including solid electrolytes.ghosal_gas_1994, mosciatti_light-modulation_2016, hou_applications_2019 They exhibit a diverse range of tunable properties dependent on both chemical structure and morphology. Copolymers, specifically block copolymers (BCPs), are a class of polymers in which multiple chemically distinct polymer chains, as building blocks, are bonded together forming one macromolecule. When the blocks of a copolymer are immiscible, unfavorable interactions between the blocks lead to phase separation resulting in unique morphologies.darling_directing_2007 These unique nanoscale phase-separated microdomains can be leveraged to facilitate the design of advanced materials.heeger_25th_2014, liu_unraveling_2018 For example, BCP membranes exhibit strong potential to surpass traditional organic and inorganic membranes by offering improved selectivity, permeability, and pore uniformity for applications such as water purification, battery electrolytes, and gas separation fields.wang_recent_2024
@@
BCP self-assembly behavior at thermodynamic equilibrium can be accurately predicted using self-consistent field theory.bates_block_1990, helfand_block_1976, leibler_theory_1980 The theory-predicted morphology of BCPs is primarily dictated by interactions between the blocks, known as the Flory-Huggins interaction parameter $\chi$}, the degree of polymerization $N$, and the volume fraction of each block $\phi$}. matsen_unifying_1996 In solid state, the phase diagram can increase in complexity due to polymer chains becoming kinetically trapped during sample preparation, resulting in a non-equilibrium state. Furthermore, the complexity increases in BCP thin films due to interfacial energy and confinement effects as the morphology now depends on the strength of interaction at interfaces.mansky_controlling_1997, han_perpendicular_2009, brassat_nanoscale_2020
@@
To achieve precise control over BCP solid state morphology, processing conditions such as solvent type, sample casting method, and post-deposition annealing technique require optimization.albalak_thermal_1997 The importance of solvent type is especially pronounced when the solvent preferentially swells one block, altering the volume fraction and thus the morphology.albalak_solvent_1998 Additionally, annealing techniques such as thermal and solvent-vapor annealing of BCP thin films play a large role in the final morphology. These techniques allow for increased mobility of the polymer chains, leading to a more thermodynamically favorable morphology.wang_recent_2024 In such cases, the kinetic effects of temperature change and solvent removal also play a key role in the final morphology.gu_situ_2014, albert_systematic_2012 Moreover, casting methods such as drop casting and spin coating of BCP solution also affect the morphology mainly due to variation in solvent evaporation dynamics. Spin coating promotes rapid solvent removal and shear induced alignment resulting in uniform films with well-ordered microdomains and reduced defect densities.zhang_influence_2012, tyona_theoritical_2013 On the other hand, slower solvent evaporation during drop casting leads to thicker films with varied domain orientation and rougher surface features.zhang_influence_2012, tyona_theoritical_2013, jung_effect_2010 Therefore, these differences affect overall phase behavior and domain spacing, as the interplay between solvent evaporation, interfacial energetics, and film thickness determine the degree of phase separation and ordering.jung_effect_2010, mishra_effect_2012, saito_mechanical_2021
@@
Scattering and microscopy are common techniques for morphological characterization of BCP films. Grazing incidence small-angle X-ray scattering (GISAXS) is an inverse-space characterization technique measuring the phase behavior of BCP films to extract the average domain spacing and degree of ordering, where the domain of a BCP refers to the aggregation of a specific block.li_alternating_2019 Atomic force microscopy (AFM) is a real-space characterization technique that can provide insights in BCP surface morphology and serve as a complementary characterization method to GISAXS.lorenzoni_assessing_2015 Moreover, periodically ordered arrangements of BCP domains, i.e., grains, can be identified and measured in AFM as well. However, traditional analysis of BCP characterization data remains time consuming, as manual analysis is often required, especially for features such as grain size from AFM images.
Despite yielding great achievements, the traditional process of BCP solid state morphology control is constrained by laborious experimentation and data analysis required from researchers for the optimization of processing conditions. Design of Experiments (DoE) has long been used to systematically reduce the experimental burden by probing a controlled subset of processing conditions and inferring trends.cao_how_2018 However, DoE typically assumes linear or low-order polynomial behavior, limiting its ability to model complex, nonlinear behavior and interactions, especially in high-dimensional or constrained experimental spaces.williamson_design_2022 Even with advanced DoE techniques such as high-dimensional DoE, these methods may still fail to capture the full complexity of response surfaces in systems with many variables and nonlinear responses.bukys_high-dimensional_2020 Furthermore, attempts to incorporate chemical identity into these frameworks by encoding reagents or materials as continuous variables, often through single-value descriptors, offer only a partial solution.wall_practical_2025 Because such representations compress inherently multidimensional chemical features into a single number, they struggle to preserve the true chemical variability, limiting the usefulness of DoE when comparing distinct chemistries.
@@
In contrast, high-throughput data analysis and machine learning (ML) algorithms have demonstrated promising results in the field of materials science, offering more flexible approaches for capturing complex, multivariate relationships and solving optimization problems.struble_prospective_2024, upadhya_automation_2021, macleod_self-driving_2020 However, these algorithms typically require large datasets for the training process, which remains scarce for polymer systems, necessitating high-throughput experiments to generate the required data.chen_polymer_2021, patra_data-driven_2022, day_navigating_2023, schuett_application_2024 Therefore, high-throughput data acquisition and analysis, combined with ML-aided data interpretation, can be leveraged to explore the processing space for controlling the solid-state morphology of BCPs.
@@
ML algorithms are extremely diverse, tailored for specific tasks and datasets. These algorithms, in the case of supervised learning, extract relationships from a labeled datasets while optimizing a defined performance metric.jordan_machine_2015 Within supervised learning, there are two primary tasks: regression and classification. The key difference between these two tasks is the output: a numerical value for regression and a categorical output for classification.castillo-boton_machine_2022 In the field of materials science, these models have provided researchers with an approach to reducing experimentation by predicting the properties of untested systems, as well as a tool to understand the relationships within complex material systems.wei_machine_2019, ma_machine-learning-assisted_2023, lu_accelerated_2018, oliynyk_high-throughput_2016 Numerous models have been applied in the materials science field to great success, including random forest (RF), support vector machine (SVM), multilayer perceptron (MLP), and extreme gradient boosting (XGB).fridgeirsdottir_multiple_2018, tao_benchmarking_2021, kim_comparison_2021, banerjee_mathematical_2022, liang_machine-learning-assisted_2021, xu_machine_2021, chen_predicting_2021, xu_accurate_2008
Within the polymer field, ML has been leveraged to decipher scattering profiles and AFM images of polymers, elucidating polymer conformational and morphological features.tung_small_2022, ding_deciphering_2025, kobayashi_machine_2023, paruchuri_machine_2024 Specifically, the Jayaraman group has pioneered the integration of ML and molecular simulations to advance the interpretation of scattering data from complex polymer solutions. Their approach enables the analysis of scattering profiles that lack conventional fitting equations, overcoming a major limitation in characterizing novel polymer architectures. beltran-villegas_computational_2019, wu_machine_2022, wessels_machine_2021Furthermore, Doerk and coworkers demonstrated an autonomous platform for the discovery of novel BCP morphologies through a directed self-assembly approach. Their work leveraged a gaussian processor to guide the direction of future scattering experiments to further study these novel BCP morphologies.doerk_autonomous_2023 These works significantly broaden the applicability of scattering experiments, offering new insights into the structure and behavior of emerging polymer systems.
@@
Despite the emerging use of ML for polymer property predictions in the solution and solid state, comparatively little work has focused on the utilization of ML to understand how processing conditions influence the solid-state morphology of BCPs – an area critical for the design of functional materials. Furthermore, generating large datasets in the field of polymer science remains a significant challenge due to long experimentation and analysis times. Many studies have circumvented this issue by compiling pre-existing data from online databases and published papers into a single dataset.ethier_predicting_2022, kim_polymer_2021, sha_machine_2021, struble_prospective_2024 However, this compiled data can be highly variable, as it comes from different laboratories, leading to inconsistencies in both the experimental setups and data analysis, as well as missed or mislabeled descriptors.
@@
Therefore, the synergistic use of high-throughput data analysis and ML methods can address the previously mentioned challenges and accelerate the rational design of polymeric materials.eyke_toward_2021, beaucage_automation_2024 In this work, we focus on a single block copolymer system, polystyrene-block-poly(ethylene oxide) (PS-b-PEO, Mn = 23-b-7 kg/mol), which forms hexagonally packed cylindrical microdomains. This well-characterized composition serves as a model to investigate how processing conditions can drive transitions in domain orientation, as well as changes in ordering and domain size. We present an approach for the high-throughput data analysis of BCP thin film GISAXS profiles and two-dimensional grain sizes measured by AFM, creating a robust dataset of morphological properties from over 200 BCP thin films. In addition, interpretable ML is leveraged to evaluate the relationship between processing parameters and BCP morphology, and offers direct morphological property predictions from processing conditions. This work serves as a proof of concept for an integrated high-throughput data analysis-ML framework for the optimization of processing conditions for polymeric materials (Fig. 1).
@@ 2.1 Experimentation and characterization
Polystyrene-block-poly(ethylene oxide) (PS-b-PEO) (Mn = 23-b-7 kg/mol), which forms a hexagonally packed cylinder morphology, was used as a model system to generate thin films using the “nonvolatile additive solvent annealing” (NVASA) approach. The details of the NVASA process can be found in our previous publication.weller_roll--roll_2019 PS-b-PEO was dissolved overnight in host solvents composed of mixtures of toluene and tetrahydrofuran (THF) at the following volume ratios, 100:0, 90:10, 80:20, 70:30, 60:40 and 50:50, to form 2 to 3 wt% stock solutions. With NVASA, it has been previously reported that a controlled amount of high boiling point additive can briefly remain in the polymer films after the primary low boiling point solvent evaporates after spin coating, thereby swelling the BCP thin film, providing segmental mobility and improving domain order. shui_rapid_2025 Therefore, for each stock solution, varied masses of a high boiling point additive: chloronaphthalene (CN), or methylnaphthalene (MN), were added to achieve additive ratios from 1.00 to 7.00. Despite the differing amounts of additive, the rate of solvent removal has been shown to remain relatively consistent.weller_roll--roll_2019 The control of additive ratio is dictated by the mass of additive relative to the mass of BCP (Equation 1).
$\text{Additive Ratio}=\frac{mass_{additive}}{mass_{BCP}}+1$ (1)
@@
Oxford Cypher Asylum AFM was used to image the domain orientation of BCP thin films. Phase and height images ($\mum ×\times 1 μ\mum$) with resolution of 256 $\times$ 256 pixels were taken in tapping mode. The tapping mode cantilevers (Tap300AI-G) with force constant of 40 N/m and resonant frequency of 300 kHz were used to scan the surface of films at a rate of 1.95 Hz. AFM raw data were collected using Igor Pro–based software, then flattened, and exported as images using Gwyddion software. These images were later processed in high-throughput fashion to extract properties of interest.
@@ 2.2.1 GISAXS
In this equation, $I(q)$ represents the scattering intensity as a function of the scattering vector $q$. The constant $A$ accounts for a baseline background offset, while the term $Cq^{-d}$ models a power law decay commonly associated with background or form factor scattering, with $C$ as the amplitude and $d$ as the decay exponent. The Gaussian peak is characterized by an amplitude $D$, a center position $b$, and a full width at half maximum (FWHM) $w$. The peak position $b$ is used to calculate the domain spacing $ds$, using the relation $ds=2\pi/b$. The FWHM, $w$, reflects the distribution of domain spacings and serves as an indicator of the degree of structural order in the sample. This fitting function enables simultaneous characterization of both periodic structural features and background scattering behavior.gu_situ_2014 An in-house Python script was used to fit the scattering data by first locating the primary peak position, then using this value as a starting parameter for the fitting equation, which the Levenberg-Marquardt optimization algorithm was used to minimize the sum of squared residuals. The optimization process was allowed a maximum of 50,000 function evaluations, i.e., the model could be evaluated up to 50,000 times with different parameter values during the search for the optimal fit. The script outputs the final fitted values for peak position and FWHM, along with the coefficient of determination ($R^{2}$) of the fitted function. This process for GISAXS analysis is both highly accurate and computationally efficient, with an average fitting $R^{2}$ of 0.99, taking less than 1 second per sample (Fig. 2).
To verify that the domain orientations observed in AFM were consistent with the bulk thin film structure, higher-order peaks were analyzed relative to the primary peak. In the scattering profiles, higher-order peaks appeared only in samples with horizontally-oriented cylindrical domains. However, the expected (11) peak at a spacing ratio of $\sqrt{2}$ was absent; instead, the (20) peak was observed at a ratio of 2. This behavior is typical for horizontally-oriented hexagonal cylinders, where the structure factor suppresses the (11) peak.bendejacq_well-ordered_2002, chu_morphologies_1995 Because a ratio of 2 is also consistent with vertically-stacked lamellae, the spacing ratio alone could not confirm the bulk morphology. To resolve this ambiguity, 2D GISAXS patterns were analyzed (Fig. S1), confirming the presence of horizontally-oriented hexagonal cylinders rather than vertical lamellae.smarsly_quantitative_2005 These patterns are discussed in detail in the SI.
Automated peak ratio detection was carried out by isolating the region between 1.5 and 2.5 times the q-position of the primary peak, in this case corresponding to the expected (20) peak for hexagonally packed cylinders (Fig. 2). This window can be adjusted to target peaks from other morphologies. The algorithm then scans the region for a secondary peak. If none is found, no output is saved. If a peak is detected, the ratio is calculated and stored, provided it falls within a user-defined range. Otherwise, the output file flags the peak as outside the expected range, indicating further analysis is needed. Using this approach, secondary peaks were correctly identified with 92% accuracy.